OpenStudy (anonymous):

Lgbariddle >:)) Using any Maths Symbols or Signs prove it... 0 0 0 = 6 1 1 1 = 6 2 2 2 = 6 3 3 3 = 6 4 4 4 = 6 5 5 5 = 6 6 6 6 = 6 7 7 7 = 6 8 8 8 = 6 9 9 9 = 6 You can Insert Everything but Numbers, And you can't change everything.

Directrix (directrix):

0 0 0 = 6 [Cos(0) + Cos(0) + Cos(0)]! = 6 Note: Cosine of zero degrees = 1.

OpenStudy (unklerhaukus):

1 1 1 = 6 [1 + 1 + 1]! = 6

Directrix (directrix):

2+2+2 = 6

OpenStudy (unklerhaukus):

3 x 3 - 3 = 6